OpenStudy (anonymous):

7n-2<=cn how to calculate the values of c and n

OpenStudy (texaschic101):

7n - 2 <= cn ..... for n 7n - 2 <= cn ---- subtract 7n from both sides -2 <= cn - 7n -2 <= n(c - 7) -2 / (c - 7) <= n or n >= - (( 2/(c-7)) 7n - 2 <= cn...for c --- divide both sides by n (7n - 2) / n <= c or c >= (7n - 2) / n I am not 100% sure on this one

OpenStudy (anonymous):

Here you go